OpenStudy (anonymous):

Use the properties of exponents to simplify write in form b^n where b and n are real numbers. [(sqrt3)^pi]^4 please show work

OpenStudy (across):

\[(\sqrt{3}^\pi)^4=3^{2\pi}\]